OpenStudy (anonymous):

First factor the trinomial as a binomial square and then use the square root method to solve. Round the solution to the nearest tenth. x2 - 14x + 49 = 3 A. {5.3, 8.7} B. {4.8, 9.2} C. {3.5, 7.5} D. {6.2, 8.3}

OpenStudy (anonymous):

x^2 - 2(x)(7) + 7^2 = (x - 7)^2

OpenStudy (anonymous):

what is the square root method?

OpenStudy (hoblos):

(x-7)2 =3 x= +/-sqrt3 +7 ==>A

OpenStudy (anonymous):

hoblos, can you explain how you did it?

OpenStudy (hoblos):

0k ;) (x-7)2 =3 => x-7 = +/- sqrt3 => x = +/-sqrt3 +7 get it

OpenStudy (anonymous):

yes, thank you.

OpenStudy (hoblos):

welcome :D